# Extracting Problem Structure with LLMs for

### v1

@@ 1 Introduction
Local search preprocessing guides CDCL solvers to faster solutions through better starting points. Balint and Manthey (2013) showed that preprocessing improves CDCL-based SAT-solving performance. Current SAT solvers like CaDiCaL (Biere, 2019; Biere et al., 2020, 2024) and CryptoMiniSat (Soos, 2020) have adopted this approach in their preprocessing phase. These tools apply basic strategies that work well for random problems but miss critical patterns in structured instances. SAT encodings of real problems contain inherited patterns from graph layouts, data connections, and domain-specific rules. The transformation to Conjunctive Normal Form (CNF) obscures these patterns. Current local search methods skip these structures in favor of general approaches. This paper addresses these limitations by introducing a framework that leverages LLMs to generate local search strategies tailored to encoding structures, enabling solvers to take advantage of these patterns for improved performance. Our research addresses three questions:
@@
    How can LLMs analyze PySAT (Ignatiev et al., 2024) code to interpret how problem structure translates to SAT clauses?
@@
Our method applies LLMs to read and interpret the PySAT code that converts structured problems to SAT. The model spots high-level constructs like graph connections, path constraints, or counting limits and then builds these insights into a specialized local search procedure. This procedure targets the encoding rather than specific instances, making it applicable across all instances encoded this way. This use of LLMs moves beyond generic applications of AI tools by integrating them directly into the SAT solving workflow. While existing research highlights using LLMs for interactive code generation, such as in GitHub Copilot (Jiang et al., 2024), our approach addresses the challenges associated with fully automated code generation for specialized tasks like SAT preprocessing. Significantly, the strategies we generate are encoding-specific but apply across problem instances, addressing a key gap in SAT-solving techniques.
@@
Code and instances are available on Zenodo (Schidler and Szeider, 2025)
@@ 2.1 SAT solving
The propositional satisfiability problem (SAT) takes a propositional formula and asks if there exists a variable assignment that makes the formula true, i.e., that satisfies the formula. Modern complete SAT solvers use Conflict-Driven Clause Learning (CDCL) (Silva and Sakallah, 1996; Marques-Silva et al., 2021; Fichte et al., 2023) and search for a satisfying assignment or proves that no such assignment exists.
@@ 2.2 Local Search Algorithms for SAT
The GSAT algorithm (Selman et al., 1992) introduced the basic local search approach for SAT. GSAT picks the variable flip that leads to the most satisfied clauses. WalkSAT (Selman et al., 1994) operates by randomly selecting an unsatisfied clause and then choosing a variable within that clause to flip. These algorithms sparked numerous variations: Novelty (McAllester et al., 1997) considers the time since a variable’s last flip when selecting moves. Adaptive Novelty+ (Hoos, 2002) automatically tunes its noise parameter during search. ProbSAT (Balint and Schöning, 2012) uses probability distributions based on make and break values to choose variables. YalSAT (Biere, 2014) combines ideas from several algorithms with additional restart strategies.
@@ 2.3 Local Search in CDCL Solvers
Modern CDCL solvers use local search during preprocessing to find promising initial assignments. CaDiCaL (Biere, 2019; Cai et al., 2022; Biere et al., 2024) implements a variant of ProbSAT in its rephasing procedure. The solver stores good assignments found during search and uses them as starting points after restarts. This hybrid approach performs particularly well on random and hard combinatorial instances by combining the systematic nature of CDCL with local search’s ability to find solutions in satisfiable regions quickly. The integration of local search into CDCL brings two main benefits: better initial assignments can guide the solver toward solutions faster, and local search patterns can inform restart strategies. However, current implementations use generic local search methods that don’t account for problem structure.
@@ 2.4 Machine Learning and SAT solving
Machine learning has enhanced SAT solving in multiple ways. Deep neural networks can learn variable selection policies that speed up SAT solving (Selsam and Bjørner, 2019). A recent line of work leverages reinforcement learning to guide SAT local search (Yolcu and Póczos, 2019). LLMs have recently shown success in algorithm generation. AlphaCode (Li et al., 2022) and CodeGen (Nijkamp et al., 2023) can generate correct implementations of algorithms. For constraint satisfaction, system StreamLLM generates streamlining constraints using LLM-calls (Voboril et al., 2024), and MCP-Solver, on the other hand, allows interactive calls to a constraint solver within an LLM chat via the Model Context Protocol (Szeider, 2024). However, using LLMs to analyze and improve algorithms is still a new direction. Related to our approach is work that uses LLMs to optimize compiler passes (Cummins et al., 2023). For our target problems, SAT encodings remain competitive with specialized algorithms.
@@ 3 Problem Statement
We consider the problem of automatically finding problem specific local search approaches that perform well in conjunction with CDCL SAT solvers. Existing local search approaches have two drawbacks. First, these algorithms are usually general purpose algorithms that do not consider, or know of, the encoding scheme that has been used to create the formula. We expect that special considerations for the encoding scheme boosts the performance of local search. Second, it has been shown empirically that local search often struggles to find satisfying assignments on its own (Li and Li, 2012; Cai and Zhang, 2021; Cai et al., 2022). As discussed in the previous section, hybrid approaches that combine local search and CDCL solvers often achieve better results on hard combinatorial instances than either paradigm on its own (Cai et al., 2022). Hence, local search methods specifically designed for hybrid approaches that are specific to the encoding scheme promise an improved performance on these hard instances.
@@
Coloring is one of Karp’s original 21 NP-complete problems (Karp, 1972). Given a graph $G=(V,E)$ and a positive integer $k$, we need to assign each vertex one of $k$ different colors such that adjacent vertices are not monochromatic. The classic encoding scheme is comparatively simple (Gelder, 2008), checking whether $G$ permits a $k$-coloring requires $O(|V|\cdot k)$ many clauses.
@@
Given a directed graph $G=(V,A)$ and a non-negative integer $k$, the goal is to find a set of vertices with cardinality at-most $k$, such that after removing these vertices, the directed graph is acyclic. There are different ways to encode this into SAT (Janota et al., 2017). We use an encoding scheme that encodes reachability, which requires $O(|V|^{3})$ many clauses. Further, additional complexity is introduced by the cardinality constraint used to limit the number of removed vertices.
@@
Given a labeled dataset with numerical features and a non-negative integer $k$, the goal is to find a decision tree with depth at-most $k$ that correctly predicts the label (or class) for the whole dataset. We use the encoding scheme by Shati et al. (2021), which is fairly complicated as it has to encode, among other things, every possible path through a tree of depth $k$.
@@ 4.1 Standardizing Encoding Schemes
There are many ways to implement encoding schemes. Since we are defining a general approach for generating local search methods, we need a standardized way of representing encoding schemes. We express our encoding schemes in Python using the popular framework PySAT (Ignatiev et al., 2018). With PySAT, one can express the encoding scheme in terms of understandable—both by a human and an LLM—programming constructs. Further, PySAT abstracts away cardinality constraints behind a single function.
@@ 5.1 Setup
We consider 257 graphs from TreewidthLib [^3] The graphs were kindly provided by Fichte et al. (2017). and DIMACS [^4] https://sites.cc.gatech.edu/dimacs10/. For each instance, we establish an upper bound using DSATUR (Brélaz, 1979) and then improve this bound for ten hours using the SAT encoding. Further, whenever we know a better upper bound from literature (Sun et al., 2021), we decrease the upper bound by one to obtain a hard instance. This results in 10 training and 38 testing instances.
@@
We follow the evaluation in (Kiesel and Schidler, 2023) and use 513 graphs from PACE 2022 (Großmann et al., 2022), ICCMA [^5] https://argumentationcompetition.org/2021/, and random graphs according to Zhou (2016). We establish the optimal solution using DAGer (Kiesel and Schidler, 2022, 2023). We discard any instance not solvable by DAGer. This results in 18 training instances and 124 test instances.
@@
We consider 69 datasets used in related work on optimal decision trees (Bessiere et al., 2009; Olson et al., 2017; Narodytska et al., 2018; Verwer and Zhang, 2019; Avellaneda, 2020; Schidler and Szeider, 2024). We run the SAT encoding on each instance for ten hours, starting with an initial bound of 10, as the encoding becomes too large with higher bounds. We decrease the bound with every solution we find. The training set consists of 9 instances and and the test set contains 32 instances.
@@ 5.5 Test Instance Performance
BDDT is the problem the LLMs struggled most with. No matter which local search function is used, the SAT solver alone can solve more instances. It is not known how many additional instances can reasonably be expected to be solvable by a SAT solver. Hence, the local search functions might perform much better in case the test instances contained more hard but solvable instances. One of the best local search functions can aid the SAT solver to solve instance objectivity in 2483 seconds, which was reported by Schidler and Szeider (2024) as not solvable within six hours. This raises our confidence in the quality of the local search functions.

### v2

@@ 1 Introduction
Local search preprocessing guides CDCL solvers to faster solutions through better starting points. ? [[?]] showed that preprocessing improves CDCL-based SAT-solving performance. Current SAT solvers like CaDiCaL [???] and CryptoMiniSat [?] have adopted this approach in their preprocessing phase. These tools apply basic strategies that work well for random problems but miss critical patterns in structured instances. SAT encodings of real problems contain inherited patterns from graph layouts, data connections, and domain-specific rules. The transformation to Conjunctive Normal Form (CNF) obscures these patterns. Current local search methods skip these structures in favor of general approaches. This paper addresses these limitations by introducing a framework that leverages LLMs to generate local search strategies tailored to encoding structures, enabling solvers to take advantage of these patterns for improved performance. Our research addresses three questions:
@@
    How can LLMs analyze PySAT [?] code to interpret how problem structure translates to SAT clauses?
@@
Our method applies LLMs to read and interpret the PySAT code that converts structured problems to SAT. The model spots high-level constructs like graph connections, path constraints, or counting limits and then builds these insights into a specialized local search procedure. This procedure targets the encoding rather than specific instances, making it applicable across all instances encoded this way. This use of LLMs moves beyond generic applications of AI tools by integrating them directly into the SAT solving workflow. While existing research highlights using LLMs for interactive code generation, such as in GitHub Copilot [?], our approach addresses the challenges associated with fully automated code generation for specialized tasks like SAT preprocessing. Significantly, the strategies we generate are encoding-specific but apply across problem instances, addressing a key gap in SAT-solving techniques.
@@
Code and instances are available on Zenodo [?]
@@ 2.1 SAT solving
The propositional satisfiability problem (SAT) takes a propositional formula and asks if there exists a variable assignment that makes the formula true, i.e., that satisfies the formula. Modern complete SAT solvers use Conflict-Driven Clause Learning (CDCL) [???] and search for a satisfying assignment or proves that no such assignment exists.
@@ 2.2 Local Search Algorithms for SAT
The GSAT algorithm [?] introduced the basic local search approach for SAT. GSAT picks the variable flip that leads to the most satisfied clauses. WalkSAT [?] operates by randomly selecting an unsatisfied clause and then choosing a variable within that clause to flip. These algorithms sparked numerous variations: Novelty [?] considers the time since a variable’s last flip when selecting moves. Adaptive Novelty+ [?] automatically tunes its noise parameter during search. ProbSAT [?] uses probability distributions based on make and break values to choose variables. YalSAT [?] combines ideas from several algorithms with additional restart strategies.
@@ 2.3 Local Search in CDCL Solvers
Modern CDCL solvers use local search during preprocessing to find promising initial assignments. CaDiCaL [???] implements a variant of ProbSAT in its rephasing procedure. The solver stores good assignments found during search and uses them as starting points after restarts. This hybrid approach performs particularly well on random and hard combinatorial instances by combining the systematic nature of CDCL with local search’s ability to find solutions in satisfiable regions quickly. The integration of local search into CDCL brings two main benefits: better initial assignments can guide the solver toward solutions faster, and local search patterns can inform restart strategies. However, current implementations use generic local search methods that don’t account for problem structure.
@@ 2.4 Machine Learning and SAT solving
Machine learning has enhanced SAT solving in multiple ways. Deep neural networks can learn variable selection policies that speed up SAT solving [?]. A recent line of work leverages graph neural networks to guide SAT local search [?]. LLMs have recently shown success in algorithm generation. AlphaCode [?] and CodeGen [?] can generate correct implementations of algorithms. For constraint satisfaction, system StreamLLM generates streamlining constraints using LLM-calls [??], and MCP-Solver, on the other hand, allows interactive calls to a constraint solver within an LLM chat via the Model Context Protocol [?]. However, using LLMs to analyze and improve algorithms is still a new direction. Related to our approach is work that uses LLMs to optimize compiler passes [?]. For our target problems, SAT encodings remain competitive with specialized algorithms.
@@ 3 Problem Statement
We consider the problem of automatically finding problem specific local search approaches that perform well in conjunction with CDCL SAT solvers. Existing local search approaches have two drawbacks. First, these algorithms are usually general purpose algorithms that do not consider, or know of, the encoding scheme that has been used to create the formula. We expect that special considerations for the encoding scheme boosts the performance of local search. Second, it has been shown empirically that local search often struggles to find satisfying assignments on its own [???]. As discussed in the previous section, hybrid approaches that combine local search and CDCL solvers often achieve better results on hard combinatorial instances than either paradigm on its own [?]. Hence, local search methods specifically designed for hybrid approaches that are specific to the encoding scheme promise an improved performance on these hard instances.
@@
Coloring is one of Karp’s original 21 NP-complete problems [?]. Given a graph $G=(V,E)$ and a positive integer $k$, we need to assign each vertex one of $k$ different colors such that adjacent vertices are not monochromatic. The classic encoding scheme is comparatively simple [?], checking whether $G$ permits a $k$-coloring requires $O(|V|\cdot k)$ many clauses.
@@
Given a directed graph $G=(V,A)$ and a non-negative integer $k$, the goal is to find a set of vertices with cardinality at-most $k$, such that after removing these vertices, the directed graph is acyclic. There are different ways to encode this into SAT [?]. We use an encoding scheme that encodes reachability, which requires $O(|V|^{3})$ many clauses. Further, additional complexity is introduced by the cardinality constraint used to limit the number of removed vertices.
@@
Given a labeled dataset with numerical features and a non-negative integer $k$, the goal is to find a decision tree with depth at-most $k$ that correctly predicts the label (or class) for the whole dataset. We use the encoding scheme by ? [[?]], which is fairly complicated as it has to encode, among other things, every possible path through a tree of depth $k$.
@@ 4.1 Standardizing Encoding Schemes
There are many ways to implement encoding schemes. Since we are defining a general approach for generating local search methods, we need a standardized way of representing encoding schemes. We express our encoding schemes in Python using the popular framework PySAT [?]. With PySAT, one can express the encoding scheme in terms of understandable—both by a human and an LLM—programming constructs. Further, PySAT abstracts away cardinality constraints behind a single function.
@@ 5.1 Setup
We consider 257 graphs from TreewidthLib [^3] The graphs were kindly provided by ? [[?]]. and DIMACS [^4] https://sites.cc.gatech.edu/dimacs10/. For each instance, we establish an upper bound using DSATUR [?] and then improve this bound for ten hours using the SAT encoding. Further, whenever we know a better upper bound from literature [?], we decrease the upper bound by one to obtain a hard instance. This results in 10 training and 38 testing instances.
@@
We follow the evaluation in [?] and use 513 graphs from PACE 2022 [?], ICCMA [^5] https://argumentationcompetition.org/2021/, and random graphs according to ? [[?]]. We establish the optimal solution using DAGer [??]. We discard any instance not solvable by DAGer. This results in 18 training instances and 124 test instances.
@@
We consider 69 datasets used in related work on optimal decision trees [??????]. We run the SAT encoding on each instance for ten hours, starting with an initial bound of 10, as the encoding becomes too large with higher bounds. We decrease the bound with every solution we find. The training set consists of 9 instances and and the test set contains 32 instances.
@@ 5.5 Test Instance Performance
BDDT is the problem the LLMs struggled most with. No matter which local search function is used, the SAT solver alone can solve more instances. It is not known how many additional instances can reasonably be expected to be solvable by a SAT solver. Hence, the local search functions might perform much better in case the test instances contained more hard but solvable instances. One of the best local search functions can aid the SAT solver to solve instance objectivity in 2483 seconds, which was reported by ? [[?]] as not solvable within six hours. This raises our confidence in the quality of the local search functions.